# Bidirectional Decoding: Improving Action Chunking via Closed-Loop Resampling

### v2

@@ C.1 Definitions
$\alpha_{f}$ $=\min_{\pi_{(c,h+d)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h+d)},\pi^{*})\Big{|}C\right]-\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$
@@
C.2 Discussion on Assumption 2
Before we prove the next result, we briefly discuss Assumption 2. Note that, using law of total probability, we can already write:
@@ C.2 Discussion on Assumption 2
Here $\hat{P}$ is the policy’s learned environment dynamics. Assumption 2 allows us to replace $\hat{P}$ with the true environment dynamics $P$. In other words, we assume that an optimal policy has already learned these distributions as optimally as possible. Given we are talking about the optimal policy trained with infinite data, this is a reasonable assumption. Note that this does not make inference trivial - the policy learns the distribution but, given the distribution has non-zero entropy, the policy can still infer the wrong state.
@@
C.3 Consistency-Reactivity Inequalities
@@
Proposition 1 (Consistency-Reactivity Inequalities - Upper Bound). Let $\mathcal{L}$ be a non-linear, convex loss function. Let $\mathcal{S}^{+}\subset\{s_{t-k:t}\}$ be the states both models observe and let $\mathcal{S}^{-}\coloneqq\{s_{t-k:t}\}\setminus\mathcal{S}^{+}$. Let $C\coloneqq\{a_{t-h-d:t-1}\}\cup\mathcal{S}^{+}$, $G\coloneqq\{a_{t},z_{t-k:t}\}\cup\mathcal{S}^{-}$. Then, we can bound the expected loss of the $(c,h+d)$-policy and the $(c,h)$-policy as:
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]$ $\leq\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right]-\alpha_{b}+\epsilon_{f}(1-P_{f}^{2d}).$
@@
For ease of notation, we will write $x_{a:}^{b}$ to mean $x_{a:b}$. Additionally, for greater clarity, we will explicitly include the context length of each model, so $\pi_{(c,h)}=\pi_{h}$ and $\pi_{(c,h+d)}=\pi_{h+d}$. We start by writing, using Assumption 1,
@@
Now, we group all the terms into two - one representing where the learner’s simulation matches the reality and one where it does not. Continuing from where we left off, first, define $P_{{\hat{s}=s}}^{f}\coloneqq P({s}_{t-h-d+1:}^{t-h}|s_{t-h-d},a_{t-h-d:}^{t-h-1})$
@@
Substituting these two terms back in, we conclude. ∎
Now, we prove the lower bound:
Proposition 1 (Consistency-Reactivity Inequalities - Lower Bound). Let $\mathcal{L}$ be a non-linear, convex loss function. Let $\mathcal{S}^{+}\subset\{s_{t-k:t}\}$ be the states both models observe and let $\mathcal{S}^{-}\coloneqq\{s_{t-k:t}\}\setminus\mathcal{S}^{+}$. Let $C\coloneqq\{a_{t-h-d:t-1}\}\cup\mathcal{S}^{+}$, $G\coloneqq\{a_{t},z_{t-k:t}\}\cup\mathcal{S}^{-}$. Then, we can bound the expected loss of the $(c,h+d)$-policy and the $(c,h)$-policy as:
$\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right]$ $\leq\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]-\alpha_{f}+\epsilon_{b}(1-P_{b}^{2d}).$
Proof.
We proceed in a manner similar to the proof of the upper bound. For ease of notation, we will write $x_{a:}^{b}$ to mean $x_{a:b}$. Additionally, for greater clarity, we will explicitly include the context length of each model, so $\pi_{(c,h)}=\pi_{h}$ and $\pi_{(c,h+d)}=\pi_{h+d}$. We start by writing, using Assumption 1,
@@
C.4 Discussion on Assumption 1
We assumed that $c+h<k$ so that the larger action chunk model can condition on more states from the distant past that are temporally correlated with $a_{t}$. In the case where $c+h\geq k$, the larger action chunk model will not get any advantage (so, $\alpha=0$) and can only suffer from not having observed the recent past states. However, this is a very unlikely scenario because we expect human demonstrators to have a large memory horizon i.e.$k$ is expected to be large.
C.5 Proof of Corollary 2 and Corollary 3
Now, we prove Corollary 2 as a direct consequence of the Consistency-Reactivity Inequalities. In a near-deterministic environment, $P_{f}$ is close to 1. This is because, conditioned on the state and action at time $t^{\prime}-1$, we can confidently infer the state visited at time $t^{\prime}$ as the environment lacks noise.
Corollary 2 (Restated) In a near-deterministic environment, if $a_{t}$ is temporally dependent on at least one state in $\{s_{t-h-c-d:t-h-c-1}\}$ and $\epsilon_{f}$ is finite,
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]<$ $\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right]$
@@ C.5 Proof of Corollary 2 and Corollary 3
This follows from Proposition 1 by taking $P_{f}\approx 1$, $\epsilon_{f}$ not large and $\alpha_{b}>0$ (since $a_{t}$ is temporally dependent on at least one state in $\{s_{t-h-c-d:t-h-c-1}\}$ and the states in $\mathcal{T}_{b}$ cannot be deterministically inferred from the context of $\pi_{h}$). ∎
Lastly, we prove Corollary 3. We assume that, in a highly stochastic environment, $P_{b}$ is small. This is because, for any time step $t^{\prime}$, the agent could reach $s_{t^{\prime}+1}$ from many different states at time $t^{\prime}$ due to the noise in the environment.
Corollary 3 (Restated) In a highly stochastic environment, if temporal dependency decreases over the number of time steps i.e. $\alpha_{f}>\epsilon_{b}$, then
@@
Starting from Proposition 1, with $P_{b}$ small (since the environment is highly stochastic), we get
$\min_{\pi_{(c,h)}^{t}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h)}^{t}-\pi_{E})|C\right]\leq\min_{\pi_{(c,h+d)}^{t}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h+d)}^{t}-\pi_{E}^{t})|C\right]-\alpha_{f}+\epsilon_{b}.$
Since temporal dependency reduces as number of steps grow, $a_{t}$ is more temporally dependent on the recent past states than on the distant past, so $\alpha_{f}>\epsilon_{b}$. With this observation, we get:
$\min_{\pi_{(c,h)}^{t}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h)}^{t}-\pi_{E})|C\right]<\min_{\pi_{(c,h+d)}^{t}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h+d)}^{t}-\pi_{E}^{t})|C\right].$
C.6 Closed-Loop versus Open-Loop in Highly Stochastic and Near-Deterministic Environments

### v3

@@ D.1 Definitions
$\allowdisplaybreaks\alpha_{f}$ $=\min_{\pi_{(c,h+d)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h+d)},\pi^{*})\Big{|}C\right]-\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$
@@
D.2 Discussion on Assumption 2
Before we prove the next result, we briefly discuss Assumption 2. Note that, using law of total probability, we can write:
@@ D.2 Discussion on Assumption 2
Here $\hat{P}$ is the policy’s learned (implicit) dynamics model. Assumption 2 allows us to replace $\hat{P}$ with the true dynamics model $P$. Given that we discuss optimal policies for each choice of action horizon trained with infinite data, this is a reasonable. Note that this does not make inference trivial - the policy learns the distribution but, given the distribution has non-zero entropy in a non-deterministic environment, the policy can still infer the wrong state.
@@
D.3 Consistency-Reactivity Inequalities
@@
Proposition 1 (Consistency-Reactivity Inequalities). Let $\mathcal{L}$ be a non-linear convex function measuring the prediction error with respect to demonstrations. Let $\mathcal{S}^{+}\subset\{s_{t-k:t}\}$ be the states both the $(c,h)$ and the $(c,h+d)$ policies observe and let $\mathcal{S}^{-}\coloneqq\{s_{t-k:t}\}\setminus\mathcal{S}^{+}$. Similarly, let $\mathcal{P}^{+}\coloneqq\{a_{t-h:t-1}\}$ and $\mathcal{P}^{-}\coloneqq\{a_{t-h-d:t-h-1}\}$. Let $C\coloneqq\mathcal{P}^{+}\cup\mathcal{S}^{+}$, $G\coloneqq\{a_{t},z_{t-k:t}\}\cup\mathcal{P}^{-}\cup\mathcal{S}^{-}$. Then, we can bound the expected loss of the $(c,h+d)$-policy and the $(c,h)$-policy as:
$\alpha_{f}-\epsilon_{b}(1-P_{b}^{2d})\leq\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]-\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right]\leq-\alpha_{b}+\epsilon_{f}(1-P_{f}^{2d}).$
@@
We first prove the upper bound. For ease of notation, we will write $x_{a:}^{b}$ to mean $x_{a:b}$. Additionally, for greater clarity, we will explicitly include the context length of each model, so $\pi_{(c,h)}=\pi_{h}$ and $\pi_{(c,h+d)}=\pi_{h+d}$. We start by writing, using Assumption 1,
@@
Now, we group all the terms into two - one representing where the learner’s simulation matches the reality and one where it does not. Continuing from where we left off, first, define $P_{{\hat{s}=s}}^{f}\coloneqq P({s}_{t-h-d+1:}^{t-h}|s_{t-h-d},a_{t-h-d:}^{t-h-1})$. Then,
@@
Substituting these two terms back in, we conclude.
Now, we prove the lower bound. We proceed in a manner similar to the proof of the upper bound. For ease of notation, we will write $x_{a:}^{b}$ to mean $x_{a:b}$. Additionally, for greater clarity, we will explicitly include the context length of each model, so $\pi_{(c,h)}=\pi_{h}$ and $\pi_{(c,h+d)}=\pi_{h+d}$. We start by writing, using Assumption 1,
@@
D.4 Discussion on Assumption 1
We assumed that $c+h<k$ so that the larger action chunk model can condition on more states from the distant past that are temporally correlated with $a_{t}$. In the case where $c+h\geq k$, the larger action chunk model will not get any advantage (thus, $\alpha=0$) and can only suffer from not having observed the recent past states. However, this is a very unlikely scenario because we expect human demonstrators to have a large memory horizon $k$.
D.5 Proof of Corollary 2 and Corollary 3
Now, we prove Corollary 2 as a direct consequence of the Consistency-Reactivity Inequalities. Recall that in a near-deterministic environment, $P_{f}$ is close to 1 as the transitions are purely determined.
Corollary 2 In a highly deterministic environment, if $a_{t}$ is temporally dependent on at least one state in $\{s_{t-h-c-d:t-h-c-1}\}$ and $\epsilon_{f}$ is finite, then
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})\mid C\right]<$ $\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})\mid C\right]$
@@ D.5 Proof of Corollary 2 and Corollary 3
This follows from the upper bound of Proposition 1. We take $P_{f}\approx 1$ to get
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})\mid C\right]-\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})\mid C\right]\leq-\alpha_{b}+\epsilon_{f}(1-P_{f}^{2d})=-\alpha_{b}$
since $\epsilon_{f}$ is finite. Since $a_{t}$ is temporally dependent on at least one state in $\{s_{t-h-c-d:t-h-c-1}\}$, by Proposition 6, we have that $\alpha_{b}>0$. Therefore,
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})\mid C\right]-\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})\mid C\right]\leq-\alpha_{b}<0.$
Note that the lower bound becomes $\alpha_{f}-\epsilon_{b}(1-P_{b}^{2d})=-\epsilon_{b}(1-P_{b}^{2d})$ as $\alpha_{f}=0$ when $P_{f}=0$ (see Proposition 7). As such, we are guaranteed that both the upper and lower bounds are negative. ∎
Next, we prove Corollary 3. Note that $P_{b}$ has higher entropy since it is not conditioned on any action. In particular, we characterize a highly stochastic environment as one where $P_{b}$ is small due to the large amount of noise.
Corollary 3 In a highly stochastic environment, if temporal dependency decreases over time, i.e., $\alpha_{f}>\epsilon_{b}$, then
@@
Starting from the lower bound of Proposition 1, since the environment is highly stochastic, we have that $P_{b}$ is small. We get:
$\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h}-\pi^{*})|C\right]\leq\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d}-\pi^{*})|C\right]-\alpha_{f}+\epsilon_{b}.$
By hypothesis, $\alpha_{f}>\epsilon_{b}$. Therefore, this reduces to:
$\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h}-\pi^{*})|C\right]<\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d}-\pi^{*})|C\right]$
Note that the upper bound from Proposition 1 becomes $-\alpha_{b}+\epsilon_{f}$ as $P_{f}\approx 0$ in a highly stochastic environment. By Proposition 8, we have that $\epsilon_{f}-\alpha_{b}>\alpha_{f}-\alpha_{b}>\alpha_{f}-\epsilon_{b}>0$. Therefore, both the bounds are positive. ∎
To sanity check the Consistency-Reactivity Inequalities, we prove the next proposition. This shows that the right-hand side of the inequalities is greater than or equal to the left-hand side.
Proposition 9.
In all environments,
$-\epsilon_{b}(1-P_{b}^{2d})\leq-\alpha_{b}$
and
$\alpha_{f}\leq\epsilon_{f}(1-P_{f}^{2d}).$
Proof.
We prove the first inequality; the proof of the second proceeds similarly. Using the definition of $\alpha_{b}$ and suppressing $G$ and $C$ (as in Proposition 1) for clarity, we have:
$\alpha_{b}$ $=-\min_{\pi_{h+d}}\mathbb{E}[\mathcal{L}(\pi_{h+d},\pi^{*})]+\min_{\pi_{h}}\mathbb{E}[\mathcal{L}(\pi_{h},\pi^{*})]$
$\leq-\min_{\pi_{h+d}}\mathbb{E}[\mathcal{L}(\pi_{h+d},\pi^{*})]$
$\quad+\min_{\pi_{h+d}}\mathbb{E}\Bigl{[}\mathcal{L}\Bigl{(}P_{\text{correct inference}}\pi_{h+d}(\text{correct inference})$
$\qquad+P_{\text{incorrect inference}}\pi_{h+d}(\text{incorrect inference}),\pi^{*}\Bigr{)}\Bigr{]}$
$\leq-\min_{\pi_{h+d}}\mathbb{E}[\mathcal{L}(\pi_{h+d},\pi^{*})]$
$\quad+\min_{\pi_{h+d}}\mathbb{E}\Bigl{[}P_{\text{correct inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{correct inference}),\pi^{*}\bigr{)}$
$\qquad+P_{\text{incorrect inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{incorrect inference}),\pi^{*}\bigr{)}\Bigr{]}$
$\leq-\min_{\pi_{h+d}}\mathbb{E}[\mathcal{L}(\pi_{h+d},\pi^{*})]$
$\quad+\min_{\pi_{h+d}}\Bigl{(}\mathbb{E}\bigl{[}P_{\text{correct inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{correct inference}),\pi^{*}\bigr{)}\bigr{]}$
$\qquad+\mathbb{E}\bigl{[}P_{\text{incorrect inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{incorrect inference}),\pi^{*}\bigr{)}\bigr{]}\Bigr{)}$
$\leq-\min_{\pi_{h+d}}\mathbb{E}[\mathcal{L}(\pi_{h+d},\pi^{*})]$
$\quad+\min_{\pi_{h+d}}\Bigl{(}\mathbb{E}\bigl{[}P_{=g_{t^{\prime}}}P_{\text{correct inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{correct inference}),\pi^{*}\bigr{)}\bigr{]}$
$\qquad+\mathbb{E}\bigl{[}P_{\neq g_{t^{\prime}}}P_{\text{correct inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{correct inference}),\pi^{*}\bigr{)}\bigr{]}$
$\qquad+\mathbb{E}\bigl{[}P_{=g_{t^{\prime}}}P_{\text{incorrect inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{incorrect inference}),\pi^{*}\bigr{)}\bigr{]}$
$\qquad+\mathbb{E}\bigl{[}P_{\neq g_{t^{\prime}}}P_{\text{incorrect inference}}\mathcal{L}\bigl{(}\pi_{h+d}(\text{incorrect inference}),\pi^{*}\bigr{)}\bigr{]}\Bigr{)}$
$\leq\epsilon_{b}(1-P_{b}^{2d}).$
D.6 Closed-Loop versus Open-Loop in Highly Stochastic and Near-Deterministic Environments